OpenStudy (anonymous):

Save me? Final exam tomorrow, need help on number 25 in this packet

OpenStudy (anonymous):

Help on number 25 please

OpenStudy (anonymous):

- Plug in points to the function -> -4 = a^0+b => b= -5 and use that for the other point 5=a^2-5 which means a= square root of 10

OpenStudy (anonymous):

oo that was fairly simple haha, thanks mate =)

OpenStudy (anonymous):

would you mind taking a look at 27 as well?

OpenStudy (anonymous):

this you'll need the arc length formula = the degree/360 (degrees of a circle) x 2 Pi r which in this case is (25/360) x 2Pi (12 inches)= 5.24 inches

OpenStudy (anonymous):

i think you should come take my exam =) lol jk

OpenStudy (anonymous):

fancy taking on number 5? i got the answers for them, but i want to see if my methods and reasoning were correct

OpenStudy (anonymous):

hmm an easier way to figure this out is to look at Sin(x) behavior on the unit circle or know when Sin is at 0 on a graph, unfortunately I won't be able to help on this one cause it's my bed time. So wish you lick in your exam I guess =)

OpenStudy (anonymous):

luck

OpenStudy (anonymous):

haha its alright, i believe i have it figured out anyway, but thanks for the help mate good night =)

OpenStudy (anonymous):

a) x = 60° b) x = 90° c) x = 45°

OpenStudy (anonymous):

would you mind showing the work as to how you got to those answers?

OpenStudy (anonymous):

a) (2sinx-sqrt3)(sinx-2) = 0 2sinx - sqrt3 = 0 2sinx = sqrt 3 sin x = sqrt 3/2 = pi/3 , 2pi/3 and sinx = 2 doesnt work right?

OpenStudy (anonymous):

given in radians

OpenStudy (anonymous):

b) 2sin^2x + sin x - 3 = 0 2sin^2x + sin = 3 sinx(2sinx+1) = 3 sinx = 3 is false 2sinx + 1 = 3 2 sinx = 2 sin x = 1 which is = to pi/2